# The Decoupling of Binaries from Their Circumbinary Disks

### v2

@@ 4.3 Gravitational Waves
Figure 5: The evolution of black hole binaries in the LISA band at $z=2$ over the final three years prior to merger. Purple, green, and blue dots mark the points where each binary reaches $a_{d}$, given by Equation 7), for $\nu_{0}=\{0.1,0.01,0.001\}$ respectively. The dashed markings along each line mark one year, one month, one day, and one hour before merger. Depending on the system mass and disk viscosity, many binaries might decouple from their disks while emitting in the LISA band. This could identify the host galaxy of the merger, facilitating multi-messenger observations.
@@
To estimate the effect of gas on the phase of GWs (e.g. Derdzinski et al., 2021; Garg et al., 2022; Zwick et al., 2023), we parameterize the rate of binary orbital evolution as $\dot{a}_{b}\approx\aleph a_{b}\dot{M}/M$, where we measure $\aleph\sim 2-3$ in our present simulations, but $\aleph$ decreases sharply with Mach number in the coupled regime (e.g. $\aleph\sim-90$ at $\mathcal{M}=50,$ Dittmann & Ryan, 2022, Section 4.3), although during decoupling $\aleph$ should increase as streams of gas in the cavity become dynamically negligible. We parameterize the accretion rate as some fraction of the Eddington-limited rate $\dot{M}=\eta\dot{M}_{\rm Edd}$, where $\dot{M}_{\rm Edd}\equiv 4\pi GMc^{-1}\epsilon^{-1}\kappa^{-1}$ where we will assume a Thomson opacity $\kappa=0.4\,{\rm cm^{2}\,g^{-1}}$ and the efficiency at which rest mass is converted to radiation $\epsilon$ to be 10%. The dynamical importance of gas on the inspiral can be characterized using the ratio
@@
[…]
@@ 5 Summary
Higher-viscosity $\nu\gtrsim 0.03\,GM/c$ disks decouple very near merger ($a_{b}\lesssim 15\,R_{g}$) and exhibit variability and accretion morphology similar to coupled binaries well into the highly-relativiscic regime. Lower viscosities $\nu\lesssim 0.01\,GM/c$, appropriate to thin ($\mathcal{M}\gtrsim 10$) disks (Shakura & Sunyaev, 1973), lead to decoupling at wider separations. In such systems the accretion rate sharply declines approaching merger and the binary leaves behind a large eccentric cavity, accreting primarily from its minidisks with negligible periodicity. We showed where in the LISA band various binaries decouple, and illustrated that while dynamically inconsequential, gas can leave a measureable imprint on the GW phase observed by LISA.

### v3

@@ IV.3 Gravitational Waves
Figure 5: The evolution of black hole binaries in the LISA band at $z=2$ over the final three years prior to merger. Purple, green, and blue dots mark the points where each binary reaches $a_{d}$, given by Equation 7), for $\nu_{0}=\{0.001,0.01,0.1\}$ respectively. The dashed markings along each line mark one year, one month, one day, and one hour before merger. Depending on the system mass and disk viscosity, many binaries might decouple from their disks while emitting in the LISA band. This could identify the host galaxy of the merger, facilitating multi-messenger observations.
@@
To estimate the effect of gas on the phase of GWs (e.g. Derdzinski et al., 2021; Garg et al., 2022; Zwick et al., 2023), we parameterize the rate of binary orbital evolution as $\dot{a}_{b}\approx\aleph a_{b}\dot{M}/M$, where we measure $\aleph\sim 2-3$ in our present simulations, but $\aleph$ decreases sharply with Mach number in the coupled regime (e.g. $\aleph\sim-90$ at $\mathcal{M}=50,$ Dittmann and Ryan, 2022, Section 4.3), although during decoupling $\aleph$ should increase as streams of gas in the cavity become dynamically negligible. We parameterize the accretion rate as some fraction of the Eddington-limited rate $\dot{M}=\eta\dot{M}_{\rm Edd}$, where $\dot{M}_{\rm Edd}\equiv 4\pi GMc^{-1}\epsilon^{-1}\kappa^{-1}$ where we will assume a Thomson opacity $\kappa=0.4\,{\rm cm^{2}\,g^{-1}}$ and the efficiency at which rest mass is converted to radiation $\epsilon$ to be 10%. The dynamical importance of gas on the inspiral can be characterized using the ratio
@@
The accumulated phase of a GW event for a circular binary is $\phi=2\pi\int_{a_{i}}^{a_{f}}f(a_{b})da_{b}/\dot{a}_{b}$ between initial and final separations $a_{i}$ and $a_{f}$. When $\dot{a}_{\rm gas}\ll\dot{a}_{\rm GW}$ the cumulative deviation in phase due to gas interactions is approximately $\delta\phi\approx 2\pi\int_{a_{i}}^{a_{f}}f(a_{b})\dot{a}_{\rm gas}/\dot{a}^{2}_{\rm GW}$. Then, because
@@ V Summary
Higher-viscosity $\nu\gtrsim 0.03\,GM/c$ disks decouple very near merger ($a_{b}\lesssim 15\,R_{g}$) and exhibit variability and accretion morphology similar to coupled binaries well into the highly-relativiscic regime. Lower viscosities $\nu\lesssim 0.01\,GM/c$, appropriate to thin ($\mathcal{M}\gtrsim 10$) disks (Shakura and Sunyaev, 1973), lead to decoupling at wider separations. In such systems the accretion rate sharply declines approaching merger and the binary leaves behind a large eccentric cavity, accreting primarily from its minidisks with negligible periodicity. We showed where in the LISA band various binaries decouple, and illustrated that while dynamically inconsequential, gas can leave a measureable imprint on the GW phase observed by LISA.